Find the 100th partial sum of 2+10+8+14+14+18

1
Expert's answer
2021-10-29T02:52:45-0400
"2+10+8+14+14+18+..."

"=(2+8+14+...)+(10+14+18+...)"

Arithmetic progression "2, 8, 14, ..."


"a=2, d=8-2=6"


"S_{50}=\dfrac{2(2)+6(50-1)}{2}\cdot50=7450"

Arithmetic progression "10, 14, 18, ..."


"a=10, d=14-10=4"


"S_{50}=\dfrac{2(10)+4(50-1)}{2}\cdot50=5400"

The 100th partial sum is

"7450+5400=12850"


Or


"2+10+8+14+14+18+..."

"=12+22+32+..."


Arithmetic progression "10, 14, 18, ..."


"a=12, d=22-12=10"


"S_{50}=\dfrac{2(12)+10(50-1)}{2}\cdot50=12850"

The 100th partial sum is "12850."